OpenStudy (sarahezzmck):

Simplify completely quantity x squared plus 4 x minus 5 over quantity 5 x squared minus 8 x plus 3 times quantity 20 x minus 12 over x squared minus 6 x minus 55 how do I simplify this?

OpenStudy (campbell_st):

well you have \[\frac{x^2 +4x - 5}{5x^2 -8x + 3} \times \frac{20x - 12}{x^2 -6x - 55}\] start by doing some factoring \[\frac{(x + 5)(x -1)}{(x -1)(5x - 3)} \times \frac{4(5x -3)}{(x - 11)(x + 5)}\] cancel common factors and then see whats left...

OpenStudy (sarahezzmck):

So you cancel (x+5) and 4(5x-3)?

OpenStudy (campbell_st):

and you can also cancel (x -1) so whats left...?

OpenStudy (campbell_st):

the 4 won't cancel... just (5x -3)

OpenStudy (sarahezzmck):

Yes, I caught that, I apologize. So the remanders are 4/(x-11)?

OpenStudy (sarahezzmck):

So the answer is 4/x-11!! OMG thank you :)

OpenStudy (campbell_st):

yep... thats it... all with a little factoring.

OpenStudy (sarahezzmck):

Quick question, idk If you know how to answer this but what happens when you have a remainder?

OpenStudy (campbell_st):

nothing.... its just just expressed as a fraction 4/(x -11) its not really a remainder... the the result of multiplying 2 fractions...

OpenStudy (sarahezzmck):

I mean, I have a DBA in a couple of minutes, the question that I have to answer is "What happens when you have a remainder?" and I don't exactly know what happens and that might sound stupid and I'm really sorry but that's the question...

OpenStudy (campbell_st):

so 10 divided by 7 is 1 and remainder 3 so the remainder tells you its just 3 sevenths of the whole.... that left... its always written over the divisor of 7... hope that helps

OpenStudy (sarahezzmck):

so when there is a remainder it basically just turns into a fraction?

OpenStudy (campbell_st):

yes thats it

OpenStudy (sarahezzmck):

Thank you! Could you help me for like two more questions please?

OpenStudy (campbell_st):

I can do 1 I'm have to go in about 10 mins

OpenStudy (sarahezzmck):

What is the restriction on the quotient of quantity 5 x squared minus 10 x minus 15 divided by quantity x squared minus 9 divided by quantity 6 x plus 12 divided by quantity x plus 2

OpenStudy (campbell_st):

ok.. so the problem is \[\frac{5x^2 -10x -15}{x^2 - 9} \div \frac{6x + 12}{x + 2}\] so a little factoring and you get remember dividing by a fraction, flip and multiply \[\frac{5(x - 3)(x + 1)}{(x -3)(x + 3} \times \frac{(x + 2)}{6(x + 2)}\] cancelling common factors you have \[\frac{5(x +1)}{6(x + 3)}\] so the restriction appears to be x cannor by -3... which gives a zero denominator, which is undefined... not sure if that was what you ment..

OpenStudy (sarahezzmck):

It was the first one, and I'm so sorry I was doing my DBA.